OpenStudy (anonymous):

Solve the equation. 4 1/5 + b = 9 3/5

OpenStudy (anonymous):

4 1/5 + b = 9 3/5 4 1/5 - 4 1/5 + b = 9 3/5 - 4 1/5 (taking the same amount from each side to keep it balanced) 0 + b = 5 2/5 (subtracting 4 1/5 from each side) b = 5 2/5

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome